OpenStudy (anonymous):

(p^2+p^3-37) / (p+8)

OpenStudy (earthcitizen):

\[p^3+p^2-37 = (p^2-7p+56)(p+8),r=-485\] You could use long division or factorize

OpenStudy (earthcitizen):